CAT Chat Minutes

February 21, 2005

 

Announcements and Follow-up
Mail Delivery – Steve Downey continued discussion of mail delivery with PFS.  PFS stated that a lab-wide decision was to limit the number of mail stops.  Other divisions have concerns about the number of drop points, which they would like to see increased, but nothing has changed.  

Keith Brister suggested that if we have a fixed number of drop points, then one possibility would be to redistributes the drop points, for example move one or more of the building 401 floor drop points out onto the floor.

Bill Ruzicka said that we would look into that possibility; we are the customer we should be able to decide the location of drop-points.

Web Pages - Mohan Ramanathan explained that the web server is to be switched today.   If you continue to have problems with the way web pages appear on the monitor, please contact the floor coordinators.

Phone System - If users are aware of problems receiving calls please call the CIS help desk at
2-4400 .   They will attempt to track down the problem.

Does the New PBX cover the LOMs?

The new PBX will cover the entire APS 400 area, including LOMs and CNM.   Additional information about the PBX can be found at: http://inside.anl.gov/resources/computer_users/cis_memos/pdfs/2005_pbx_upgrade.pdf

Users are not receiving the broadcast email messages from APS auto-info (in this case about PBX problems).

Kevin spoke with Steve Leatherman and Ken Sidorowicz about setting up a list server to add to the various auto-info messages – in this case BLD info.   At this point it was felt that a list server can easily get out-of-date and maintenance could be an issue (people leave/change email address, etc.).   One solution could be to mail auto-info messages to one person at each LOM (for example, secretary) and have them forward the most relevant messages to the group.   Send Kevin Randall email if you would like to be included in the plan to receive announcements.

Machine Stats
Delivered beam:   434.87 hrs.
Percentage of scheduled time:   98.60%
Downtime during period:   6.18 hrs.
Mean time between faults:   72.48 hrs.

Fault Listing of Statistics for Run 1-2005:

Fill

Reason

Length of Downtime

# 10   02/16 08:01   To   02/21 09:04   121.06  

 Fill Ongoing             

  0.01

Beamline Data
Beamtime downtime summary by APS groups for period 02/14-21/2005

DIN

Date

Beamline

Status

Duration

Group

Description

Resolution

500

02/17/05

22-ID

Closed

2 hr 40 min

AOD-DIA

BL personnel discovered their beam had shifted and were not sure if the cause was due to steering at 23-ID last night or due to recalibration of BPMs during Machine Studies on 2/14 & 2/15.

ASD-DIA is reviewing the process for saving BPM data to the database. Pursuit of hard X-Ray BPM development is under consideration.

501

02/19/05

2-ID

Closed

33 min

AOD-IT

S. Vogt at 2-ID-E reported loss of access to PV's, the " caget" and "caput" commands are not working.

S. Vogt reported that the PV's were working. ASD-IT and M. Smith looking into why this occurred; Smith believes that the gateway may have rebooted itself.

Safety
The battery recycling program is being resumed.  On Monday, February 21st, a memo was sent out via e-mail:   Due to a change in laboratory policy concerning mail service, the mail service will no longer accept the "Green Bags" previously used to recycle batteries.   Instead, Battery Collection Containers have been distributed to act as drop points for acceptable batteries on the APS site.   The following link can be used to view the . pdf file containing further information:   http://www2.aps.anl.gov/info/BLDannounce/2005/msg00025.html

A memo has been released for non-resident users with regard to electrical safety.   ESAF has been modified to denote training. A one-page guidance sheet is to be given to the non-resident user if there is electrical work. Guidance is more specific on the memo and electrical safety update.